OpenStudy (anonymous):

long division (3x to the 4th +2x squared -12) (x squared +2)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

first write the expression the one u can't get like x raised to third....use it as 0x^3 so its actual value is 0

OpenStudy (anonymous):

do you understand it or do you need some help

OpenStudy (zehanz):

Write it like this: x²+2 / 3x^4 +2x² -12 \ ... (the answer will appear on the dots). Then ask yourself how to multiply x² (the first term of x²+2) with a number to get 3x^4. This is 3x². Now multiply x²+2 with 3x² and put the result under 3x^4 + 2x² -12. Subtract. Put the 3x² on the dots and drop the -12: x²+2 / 3x^4 +2x² -12 \ 3x² 3x^4 +6x² - --------- -4x² - 12 Now do everything the same: how to multiply x² to get -4x²? By -4: x²+2 / 3x^4 +2x² -12 \ 3x² - 4 3x^4 +6x² - --------- -4x² - 12 -4x² - 8 - -------- -4 (remember: -12 - -8 = -12 +8=-4) Now you can stop. The remainder is -4. So (3x^4 +2x² -12)/(x²+2) = 3x²-4 -4/(x²+2).

OpenStudy (anonymous):

u put a lot of effort in typing those stuff...heres the medal

OpenStudy (zehanz):

Thanks pal!

OpenStudy (anonymous):

thanks so much @ZeHanz that helps alot

OpenStudy (zehanz):

YW!